OpenStudy (anonymous):

Will fan and medal Which of the following are parts of a cone? vertex lateral face lateral edges base slant height

OpenStudy (anonymous):

Lateral faces, vertex, slant height, I think... But you might want someone else to verify.

OpenStudy (anonymous):

An Indian tepee is a conical tent. Find the number of cubic feet of air in a teepee 10 ft. in diameter and 12 ft. high. 100 cu. ft. 300 cu. ft. 400 cu. ft. \[400 \pi cu. ft\]

OpenStudy (anonymous):

Your answer is, 100 cu. ft. DO you need anything else?

OpenStudy (anonymous):

If you do tag my name afterwards. @CMFink00

OpenStudy (anonymous):

Find the missing measure for a right circular cone given the following information. Find L.A. if r = 6 and h = 8 48pi 60pi 96pi

OpenStudy (anonymous):

volume = 1/3 * pi *r^2*h = 96pi

OpenStudy (anonymous):

Your answer is 96pi.

OpenStudy (anonymous):

Find the missing measure for a right circular cone given the following information. Find r if T.A. = 12pi and L.A. = 8pi. 2 4 16

OpenStudy (anonymous):

I'm going to assume TA is short for "total area" (more commonly known simply as the surface area of the cone) and LA is short for "lateral area"... where lateral area is the surface area of the cone minus the area of the circular area on the end? TA = area of circle + LA 12pi = area of circle + 8pi 4pi = area of circle Not sure what the "missing" measure is you're looking for here... but since 4pi isn't one of the possible answers, I'm going to assume you're looking for the radius r. area of circle = 4pi (pi)(r^2) = 4pi r^2 = 4 r = 2 Answer a.

OpenStudy (anonymous):

Find the slant height for a right circular cone with a radius of 3 and a height of 5. 4 √(31) √(34)

OpenStudy (anonymous):

WHat are all of the ? marks for?

OpenStudy (anonymous):

4 \[\sqrt{31}\] \[\sqrt{34}\]

OpenStudy (anonymous):

@CMFink00

OpenStudy (anonymous):

OKay sorry give me a sec.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

l=r2+h2=32+52=5.83095

OpenStudy (anonymous):

So your answer is, 5.83

OpenStudy (anonymous):

Okay well I have school so bye. I can help later. Just tag me if you need me.

OpenStudy (anonymous):

A cone with a radius of 3" has a total area of 24pi sq. in. Find the volume of the cone. 24pi 15pi 12pi

OpenStudy (anonymous):

@CMFink00

OpenStudy (anonymous):

Give me a minute.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

12pi is your answer. Anything more? If not I need to finish my homework.

OpenStudy (anonymous):

yes i have a few more

OpenStudy (anonymous):

A large dump truck will empty its full load of stone onto a small plot of land for storage. The rock forms a cone that is 9 feet tall. The truck holds 144 cubic feet of stone. What is the area of the circular base of the pile of stones? 48 square feet 48/π□feet 16 square feet 48/π□feet−−−−−−−−−√

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Please re-write your options.

OpenStudy (anonymous):

A large dump truck will empty its full load of stone onto a small plot of land for storage. The rock forms a cone that is 9 feet tall. The truck holds 144 cubic feet of stone. What is the area of the circular base of the pile of stones? 48 square feet 48/pi square feet 16 square feet \[\sqrt{48/}\] square feet

OpenStudy (anonymous):

Okay thank you give me a sec.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Okay your correct answer is, 48/pi square feet. So the second one.

OpenStudy (anonymous):

A solid gold cube with edge of 2" is melted down and recast as a cone with a 3" radius base. Find the height of the cone. 8/(9pi) 2 8/(3pi)

OpenStudy (anonymous):

In general, if one object is re-shaped into another object, both objects will have the same volume. If you find the volume of the cube, you will have the volume of the cone. Use the formula for Volume of a Cone to find its height. Sorry but I have to finish my homework. Bye tag me later if you still need help.

OpenStudy (anonymous):

Would you like the answer or do you think you can figure it out?

OpenStudy (anonymous):

I have the answer.

OpenStudy (anonymous):

the answer please im not very good at math

OpenStudy (anonymous):

i have three more after this one

OpenStudy (anonymous):

8/3pi is your answer. Why don't you work out some problems on your own, and I will be back in a few minutes. My homework won't take long and I will stay on here, I just might not chat unless you really need help.

OpenStudy (anonymous):

Find the missing measure for a right circular cone given the following information. Find h if r = 5 and V = 100pi. 4 12 30

OpenStudy (anonymous):

Look up the formula and try to figure it out. Like I said I will be back.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@YanaSidlinskiy